OpenStudy (anonymous):

Use the following three steps to find and simplify the difference quotient of the function f(x)=\frac{1}{x-6}.

OpenStudy (anonymous):

f(x) = 1/x-6

OpenStudy (anonymous):

I believe I do the correct work but for the last two steps it says I am wrong.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Hero

OpenStudy (jhannybean):

The difference quotient: \(\dfrac{f(x+h)-f(x)}{h}\) So let's begin by finding \(f(x+h)\) first. Wherever you see an x, replace it with \(x+h\)

OpenStudy (anonymous):

1/x-6 = 1/(x+h)-6

OpenStudy (anonymous):

then i'm suppose to subtract it by f(x)

OpenStudy (jhannybean):

Alright, now we plug it in to the formula. And you are already given \(f(x)\).

OpenStudy (anonymous):

1/(x+h)-6 - 1/x = -h-6/x(x+h)-6

OpenStudy (jhannybean):

\[\frac{f(x+h)-f(x)}{h} = \frac{\dfrac{1}{x+h-6}-\dfrac{1}{x-6}}{h}\]

OpenStudy (jhannybean):

When I find a common denominator, I multiply both the denominators of the fraction together.

OpenStudy (jhannybean):

\[=\frac{\dfrac{x-6 -(x+h-6)}{(x-6)(x+h-6)}}{h}\]

OpenStudy (jhannybean):

Now let's simplify both the numerator and denominator.

OpenStudy (anonymous):

isn't the top suppose to be -h

OpenStudy (anonymous):

-h/(x+h-6)(x-6)

OpenStudy (anonymous):

then -h/(x+h-6)(x-6)/h

OpenStudy (jhannybean):

Yep, you're on track. That's what I got too. \[=\frac{\dfrac{x-6-x-h+6}{(x-6)(x+h-6)}}{h} = -\frac{\dfrac{h}{(x-6)(x+h-6)}}{h}\]

OpenStudy (jhannybean):

Now remember the rule of dividing fractions? \(\dfrac{\dfrac{a}{b}}{c} \iff \dfrac{a}{bc}\)

OpenStudy (anonymous):

its -h^2/(x+h-6)(x-6)

OpenStudy (jhannybean):

We're going to apply this rule to our simplified function. We multiply \((x-6)(x+h-6)\) with \(h\).

OpenStudy (jhannybean):

Not exactly.

OpenStudy (jhannybean):

\[ -\frac{\dfrac{h}{(x-6)(x+h-6)}}{h} = -\frac{h}{h(x-6)(x+h-6)}\]Therefore both the h's cancel out.

OpenStudy (jhannybean):

\[-\frac{\cancel{h}}{\cancel{h}(x-6)(x+h-6)}\]

OpenStudy (anonymous):

ahhh i understand

OpenStudy (anonymous):

it says the bottom is not correct

OpenStudy (anonymous):

nevermind

OpenStudy (jhannybean):

when you think of a fraction in the form \(\dfrac{\dfrac{a}{b}}{c}\) think of \(\dfrac{a}{b} \cdot \dfrac{1}{c} = \dfrac{a}{bc} \)

OpenStudy (anonymous):

i forgot the - lol

OpenStudy (jhannybean):

Ahh, good good.